OpenStudy (anonymous):

PLEASE HELPP (: If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x-6y+3z=0 2x-5y+8z=7 x+7y-4z=13

OpenStudy (phantom):

plug in x= 6y-3z in eqns 2 and 3 in place of x to get 2 eqns in y and z only

OpenStudy (anonymous):

x-6y+3z=0..........................1 2x-5y+8z=7........................2 x+7y-4z=13.........................3 Take eq 1 x= 6y-3z sub this in equation 2 and 3 2( 6y-3z)-5y+8z=7..................2 6y-3z +7y-4z=13.....................3 simplify 12y-6z-5y+8z=7 then it be 7y+2z=7................2 multiplying by 7 13y-7z=13..............3 multiplying by 2 49y+14z=49 .............2 26y-14z=26...............3 adding both equations to each other to get 75 y = 75 then y=1 sub value of y in 2 or 3 to get 7+2z=7 2z=0 z= zero sub z= zero , y =1 , in eq 1 x= 6y-3z x=6 Hope I'm right . and You are welcome .. regards